# Euclid Quick Data Release (Q1)

### v1

@@ 2.3 Reliability of the redshift estimation
Associated with each solution, the spe_z_prob parameter gives the value of the integral of the PDF under the corresponding peak, over a ± $\pm\,3\,\sigma$ interval around the central value, where $\sigma$ is the width of the Gaussian fit to the peak. This provides an indicator of the reliability of the redshift estimate: a spectrum with two or more strong coherent emission lines will typically yield a PDF with a single peak whose integral is close to unity. Conversely, a spectrum such as the one in Fig. 2 will typically induce a PDF with multiple peaks, each corresponding to different potential identifications of the line. In this case, the integral of the PDF is shared among the peaks, leading to lower associated probabilities. Finally, for a spectrum without any prominent feature, the PDF will contain many faint peaks corresponding mostly to noise fluctuations, and the probability associated with each solution will be low.
The results on the success rate of the observed samples (described in Sect. 6) show that the spe_z_prob parameter is highly nonlinear and that values below about 0.99 correspond to unreliable solutions. Improvements of the reliability of the redshift estimations are under development using machine-learning algorithms that will evaluate the whole shape of the PDF rather than a single peak. However, training of these algorithms requires a large sample of objects with secure redshifts, which will be provided by measurements on the EDS. An alternative will be to use simulated data, provided that they are sufficiently representative of the real spectra.
@@ 3.1 Comparison with external redshift measurements
In future Euclid public releases, the completeness and purity of galaxy samples for cosmological investigations will be quantified using the EDS as a reference. This will provide us with an internal, fully self-consistent control sample, to quantify redshift measurement systematics that can affect galaxy clustering and the resulting inferred cosmological parameters. Lacking such an internal reference and given the broader nature of the Q1 data set, to test the quality and limitations of the released data we need here to revert to external measurements. The only appropriate data set, in terms of area and redshift coverage, is provided by the Dark Energy Spectroscopic Instrument (DESI) Early Data Release catalogue (DESI Collaboration: Adame et al. 2024), which partly covers the EDF-N.
@@
Within the DESI catalogue, we first selected objects with a reliable redshift, corresponding to: (1) quality flag z_warn$\,=0$ (no issue identified in the redshift measurement); and (2) Delta_chi2 $>45$, which ensures that there is a significant difference between the main solution and the second one. The resulting sample was then matched to Euclid objects featuring three or more spectroscopic dithers and a reliable detection in the MER catalogue, using a matching radius of $0.2″$. This produced a sample of 25 469 matched objects. In the DESI catalogue, 10 885 of these are identified as stars, 985 as quasars, and 13 599 as galaxies.
Figures 3 and 4 compare the DESI and SPE redshift measurements, for various quality selections. Figure 3 gives a first view of the quality of a Euclid ‘cosmological sample’ of $\mathrm{H}\,\alpha$ emitting galaxies, selected within the reference flux and redshift limits. This plot directly proves the current ability of SPE to recover redshifts for those galaxies that are the targets for which Euclid NISP was built. A more in-depth discussion of this selection is presented in Sect. 6.2.
@@
In the remainder of this section, we focus on the two panels of Fig. 4, which deliberately look at the data for any kind of object, varying the quality thresholds. The scope here is to test the ability of Euclid and the SPE pipeline in particular to work out of their ‘comfort zone’ and recover information beyond the cosmological goals NISP was designed for. The matched DESI sample includes in fact also galaxies outside the $\mathrm{H}\,\alpha$ visibility range, for which the spectral characterisation has to be based on emission lines other than $\mathrm{H}\,\alpha$ or on the continuum; there are also stars, for which some information can also be present in terms of absorption lines.
@@
In the upper panel of Fig. 4, the effects of the previously discussed $\mathrm{H}\,\alpha$ prior are easily noticeable as an overdensity of points in the $1.80.9<z_{\rm SPE}<$ 1.8 region. Here, no quality criteria are applied, in order to give a full illustration of the various artefacts that can affect the performance of PF SPE, in particular when dealing with measurements outside the target redshift range.
The bottom panel displays the same information, but only for galaxies with spe_z_proba0.99absent0.99\,$\geq$ 0.99≥ 0.99, showing the relative efficiency of this selection in excluding incorrect solutions. It must be noted as well that, even if the percentage of correct solutions is low in the $z\,=\,$0 – 0.9 interval (see Sect. 6 for detailed numbers), about 15% of galaxies measured by DESI in this redshift interval have a redshift measured by SPE in the same interval with spe_z_proba0.99absent0.99\,$\geq$ 0.99≥ 0.99. Considering that the redshifts measured by SPE cannot rely on any strong emission lines except for specific categories (such as AGN), these measurements are solely based on a small portion of the continuum. This is nevertheless interesting, since NISP and the EWS were designed to measure redshifts for $\mathrm{H}\,\alpha$ emitting galaxies at $z>0.9$, and limited performance is expected below this redshift.
@@
[…]
@@ 3.2 Redshift measurement uncertainties
To evaluate the precision of Euclid redshifts, we computed the relative redshift difference $(z_{\mathrm{SPE}}-z_{\mathrm{DESI}})/(1+z_{\mathrm{DESI}})$ between the redshift measured by SPE, for objects classified as galaxies and the one provided by DESI. After excluding outliers, defined as those having an absolute difference larger than 0.005, we obtain a distribution with a median value of $0.6\times 10^{-5}$ and a $1\,\sigma$ dispersion of $1.3\times 10^{-3}$, which is plotted in Fig. 5. This provides an upper limit on the Euclid redshift uncertainties, since this quantity also contains the DESI uncertainties. However, with the DESI resolution ranging between 2000 and 5500, i.e., 4–5 times better than NISP, Euclid errors should dominate this comparison. If we restrict the sample to galaxies within the Euclid target range ($1.80.9<z<$ 1.8, but no flux restriction), to ensure that at least the $\mathrm{H}\,\alpha$ emission line is used for the redshift determination if it is present, the RMS error drops to $7\times 10^{-4}$, with a median value of $2.4\times 10^{-5}$. If we additionally cut the sample at the nominal $\mathrm{H}\,\alpha$ flux limit of $2\times 10^{-16}\,\mathrm{erg\,s}^{-1}\,\mathrm{cm}^{-2}$, the dispersion further reduces to $4.9\times 10^{-4}$. Given the low resolution of NISP and the sizeable spectroscopic undersampling, this remarkable agreement confirms the quality of the wavelength calibration, as described in Euclid Collaboration: Copin et al. (2025).
Figure 5: Normalised distribution of $(z_{\rm SPE}-z_{\rm DESI})/(1+z_{\rm DESI})$ for all the objects classified by SPE as galaxies (blue) and only in the target redshift range ($1.80.9<z<$ 1.8, orange).
@@ 3.3 Types of redshift failure
As shown in Fig. 4, a significant fraction of galaxies have an incorrect redshift if only basic quality criteria are applied, and are thus not located on the one-to-one relation. There are two types of incorrect redshift visible in this figure.
A first type, which in the forthcoming Risso et al. (in prep.) paper are called ‘line interlopers’, is caused by the misidentification of a detected line. As already explained in Sect. 2.2, the SPE pipeline can attribute the wrong line to a feature. This is particularly the case if only one line can be seen in the Euclid spectral range. In this situation, the relation between the true redshift $z_{\rm true}$ and wrong redshift $z_{\rm misID}$ is
@@
If, as in the bottom panel, we select high-quality SPE galaxies (defined as having spe_z_proba $\geq\,0.99$ and a line width – expressed in terms of velocity dispersion – smaller than 680 km s-1 to exclude spurious broad features), we are left with three such features indicating line misidentifications. The one above the one-to-one diagonal corresponds to $(1+z_{\rm SPE})/(1+z_{\rm DESI})\simeq 1.76$: this results from cases where SPE has interpreted a line as [O ii] when in fact it is $\mathrm{H}\,\alpha$. We can also notice two discrete features below the one-to-one relation, corresponding to $\mathrm{H}\,\alpha$ being misidentified as $\mathrm{Pa}\,\alpha$, $(1+z_{\rm SPE})/(1+z_{\rm DESI})\simeq 0.51$, or [S iii]9530, $(1+z_{\rm SPE})/(1+z_{\rm DESI})\simeq 0.68$. Finally, inspection of the $(1+z_{\rm SPE})/(1+z_{\rm DESI})$ histogram reveals the presence of a small peak at 1.34, corresponding to $\mathrm{H}\,\alpha$ being misidentified as $\mathrm{H}\,\beta$, although we surprisingly do not find the peak corresponding to [O iii].
As can be seen by comparing with Fig. 3, such misidentifications are largely suppressed by the $\mathrm{H}\,\alpha$ prior, and this aspect will be revisited in future versions of the pipeline. We remark that no obvious misidentification is observed in the target redshift range for cosmology, $z=0.9$–1.8.
The second type of error (giving rise to what we call ‘noise interlopers’ in Risso et al. 2025) is caused by noise features or pipeline artefacts, such as unflagged residuals of the zeroth order signal, which are mistaken for real lines. In such cases, the measured redshift values are spread over a broad range. However, they are less likely to happen for redshift ranges where two real lines fall within the nominal NISP wavelength range, since this would require two spurious peaks separated by the same exact wavelength difference of two emission lines (this is the case in the range $1.81.4<z<$ 1.8 when considering [O iii]5007 and $\mathrm{H}\,\alpha$). Conversely, the $\mathrm{H}\,\alpha$ prior tends to push these spurious redshifts into the cosmology target range. As a result, the distribution of these incorrect redshifts is not uniform, as shown in Fig. 4.
@@ 4 Spectral classification
We consider two different cases. For the first, we select high-quality spectra with three or more dithers on average (see Sect. 2.1). For the second case, we additionally require a class probability higher than 0.99. The results are not significantly better for the second case. This suggests that the spectral quality and the SPE classification probability criteria alone cannot yet efficiently select objects with a good classification.
The results are satisfactory for galaxies, with around 80% success. The results for quasars are poor, with about 80% of them being classified as galaxies. The current version of the pipeline uses a single template for quasars based on the average spectrum of the SDSS quasars. It thus allows too little flexibility to fit the Euclid spectra reliably. In addition, this template has broad lines, and type II observed AGN tend to be fitted better by the galaxy templates, for example when the line-spread function is even slightly overestimated.
Only about 50% of the stars are correctly classified, while the remaining 50% leak into the galaxy class. Only 23% of these false galaxies are at $z<0.1$, and 21% of them are in the target redshift range for cosmology ($1.80.9<z_{\rm SPE}<$ 1.8). This emphasises the fact that additional non-SPE criteria will be necessary to minimise the contamination of the cosmological sample by stars, and the high quality imaging of the Euclid space telescope will undoubtedly be efficient in doing this. In contrast, about 75% of the objects classified as stars by SPE are DESI stars. The selection of star samples is thus very promising, and the combination with other quality criteria from the Euclid photometric segment should provide high-quality samples.
Table 1: Confusion matrices of the OU-SPE classification for various selections. The DESI survey is used as a reference.
@@ 5 Line-flux measurement products and performance
Once a redshift solution has been estimated for the galaxy model, the fluxes of all emission and absorption lines that are expected to be present in the spectrum are measured. It must be noted that for this Q1 release, even if an object is classified as a quasar or a star, the lines are identified at the redshift of the galaxy solution and therefore should not be trusted; this will be modified in future releases. Line fluxes are measured with two methods: ‘Direct Integration’ provides fluxes integrating the values of all the pixels of the emission line above the continuum; and ‘Gaussian Fit’ derives them from the fit of the line using a single, double, or multiple Gaussian model, depending on the line considered. Given the low spectral resolution of NISP, the $\mathrm{H}\,\alpha$ and [N ii] doublet lines Direct-Integration fluxes are measured jointly over a unique wavelength interval to produce a single flux measurement, while the Gaussian Fit provides fluxes separately for the $\mathrm{H}\,\alpha$ and [N ii] doublet. Examples of the fitting procedure results are displayed in Fig. 6 for a high S/N emission line, as well as for one close to the flux detection limit.
Figure 6: Illustrative examples of the line flux measurement performed in SPE. Two spectra are shown from the EWS at $z\simeq 1.7$–1.8, with fluxes $F(\mathrm{H}\,\alpha)\simeq 2\times 10^{-16}$ and $5.9\times 10^{-16}\,\mathrm{erg\,s}^{-1}\,\mathrm{cm}^{-2}$ (upper and lower panels, respectively). The left plots show the entire spectrum, highlighting the position of the main emission lines, and the right plots the results of the multiple Gaussian fit to the $\mathrm{H}\,\alpha$+[N ii] complex showing in different colours the contribution of the $\mathrm{H}\,\alpha$ line (in blue), of the [N ii] doublet (in orange), and their combination (in red).
For each line and each method, SPE provides the flux, the S/N, the central wavelength and the full width at half maximum. All measurements relative to emission lines are stored in the spectro_line_features_catalog_spe_line_features_cat table: for each object, this contains all the detected lines for each redshift solution. A given line can be identified either by its name (e.g., Halpha) or a numerical identifier, and the link between line names and identifiers is given in the spectro_model_catalog_spe_lines_catalog table. Therefore, to retrieve the flux of the $\mathrm{H}\,\alpha$ line for the main redshift solution of all objects, the selection should include spe_line_name=‘Halpha’ and spe_rank=0, together with any additional conditions.

### v2

@@ 2.3 Reliability of the redshift estimation
Associated with each solution, the spe_z_prob parameter gives the value of the integral of the PDF under the corresponding peak, over a ± $\pm\,3\,\sigma$ interval around the central value, where $\sigma$ is the width of the Gaussian fit to the peak. This provides an indicator of the reliability of the redshift estimate: a spectrum with two or more strong coherent emission lines will typically yield a PDF with a single peak whose integral is close to unity. Conversely, a spectrum such as the one in Fig. 2 will typically induce a PDF with multiple peaks, each corresponding to different potential identifications of the line. In this case, the integral of the PDF is shared among the peaks, leading to lower associated probabilities. Finally, for a spectrum without any prominent feature, the PDF will contain many faint peaks corresponding mostly to noise fluctuations and the probability associated with each solution will be low.
The results for the success rate of the observed samples (described in Sect. 6) show that the spe_z_prob parameter is highly non-linear and that values below about 0.99 correspond to unreliable solutions. Improvements in the reliability of the redshift estimations are under development using machine-learning algorithms that will evaluate the whole shape of the PDF rather than a single peak. However, training of these algorithms requires a large sample of objects with secure redshifts, which will be provided by measurements on the EDS. An alternative will be to use simulated data, provided that they are sufficiently representative of the real spectra.
@@ 3.1 Comparison with external redshift measurements
In future Euclid public releases, the completeness and purity of galaxy samples for cosmological investigations will be quantified using the Euclid Deep Survey (EDS) as a reference. This will provide us with an internal, fully self-consistent control sample, to quantify redshift measurement systematics that can affect galaxy clustering and the resulting inferred cosmological parameters. Lacking such an internal reference and given the broader nature of the Q1 dataset, to test the quality and limitations of the released data, we need here to resort to external measurements. The only appropriate dataset, in terms of area and redshift coverage, is provided by the Dark Energy Spectroscopic Instrument (DESI) Early Data Release catalogue (DESI Collaboration: Adame et al., 2024), which partly covers the EDF-N.
@@
Within the DESI catalogue, we first selected objects with a reliable redshift, corresponding to: (1) quality flag z_warn=0\,=0 (no issue identified in the redshift measurement); and (2) Delta_chi2 >45>45, which ensures that there is a significant difference between the main solution and the second one. The resulting sample was then matched to Euclid objects featuring three or more spectroscopic dithers and a reliable detection in the MER catalogue, using a matching radius of 0.2. This produced a sample of 25 469 matched objects. In the DESI catalogue, 10 885 of these are identified as stars, 985 as quasars, and 13 599 as galaxies.
[…]
@@
In the remainder of this section, we focus on the two panels of Fig. 4, which deliberately look at the data for any kind of object, varying the quality thresholds. The scope here is to test the ability of Euclid and the SPE pipeline in particular to work out of their ‘comfort zone’ and recover information beyond the cosmological goals NISP was designed for. The matched DESI sample in fact also includes galaxies outside the $\mathrm{H}\,\alpha$ visibility range, for which the spectral characterisation has to be based on emission lines other than $\mathrm{H}\,\alpha$ or on the continuum; there are also stars, for which some information can also be present in terms of absorption lines.
@@
In the upper panel of Fig. 4, the effects of the previously discussed $\mathrm{H}\,\alpha$ prior are easily noticeable as an over-density of points in the $0.9<z_{\rm SPE}<1.8$ region. Here, no quality criteria are applied, in order to give a full illustration of the various artefacts that can affect the performance of the SPE PF, in particular when dealing with measurements outside the target redshift range.
[…]
@@
Figure 4: Comparison of SPE redshift measurements with those for all objects in common (including non-emitting galaxies and stars) from the DESI Early Data Release. Top: All Q1 objects with at least three observed Euclid dithers. This panel deliberately does not include any quality selection, so as to evidence the impact of the $\mathrm{H}\,\alpha$ prior (Sect. 2.2 and 3.1) and the limitations of spectroscopic data in the Q1 release beyond the Euclid cosmology selection of Fig. 3 (see Sect. 3.3). Bottom: Same as top panel, but now selecting only high-quality objects classified as galaxies (i.e. with a SPE redshift probability >0.99>0.99 and line width strictly narrower than the prior limit of $680\,$\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$). The colour coding indicates the local density of points, with black corresponding to the highest density.
@@ 3.2 Redshift measurement uncertainties
To evaluate the precision of Euclid redshifts, we computed the relative redshift difference $(z_{\mathrm{SPE}}-z_{\mathrm{DESI}})/(1+z_{\mathrm{DESI}})$ between the redshift measured by SPE, for objects classified as galaxies and the one provided by DESI. After excluding outliers, defined as those having an absolute difference larger than 0.005, we obtain a distribution with a median value of $0.6\times 10^{-5}$ and a $1\,\sigma$ dispersion of $1.3\times 10^{-3}$, which is plotted in Fig. 5. This provides an upper limit on the Euclid redshift uncertainties, since this quantity also contains the DESI uncertainties. However, with the DESI resolution ranging between 2000 and 5500, that is 4–5 times better than NISP, Euclid errors should dominate this comparison. If we restrict the sample to galaxies within the Euclid target range ($0.9<z<1.8$, but without flux restriction), to ensure that at least the $\mathrm{H}\,\alpha$ emission line is used for the redshift determination if it is present, the RMS error drops to $7\times 10^{-4}$, with a median value of $2.4\times 10^{-5}$. If we additionally cut the sample at the nominal $\mathrm{H}\,\alpha$ flux limit of $2\times 10^{-16}\,\mathrm{erg\,s}^{-1}\,\mathrm{cm}^{-2}$, the dispersion further reduces to $4.9\times 10^{-4}$. Given the low resolution of NISP and the sizeable spectroscopic undersampling, this remarkable agreement confirms the quality of the wavelength calibration, as is described in Euclid Collaboration: Copin et al. (2025).
Figure 5: Normalised distribution of $(z_{\rm SPE}-z_{\rm DESI})/(1+z_{\rm DESI})$ for all the objects classified by SPE as galaxies (blue) and only in the target redshift range ($0.9<z<1.8$, orange).
@@ 3.3 Types of redshift failure
As is shown in Fig. 4, a significant fraction of galaxies have an incorrect redshift if only basic quality criteria are applied, and are thus not located on the one-to-one relation. There are two types of incorrect redshift visible in this figure.
A first type, which in the forthcoming Risso et al. (2025) paper are called ‘line interlopers’, is caused by the misidentification of a detected line. As was already explained in Sect. 2.2, the SPE pipeline can attribute the wrong line to a feature. This is particularly the case if only one line can be seen in the Euclid spectral range. In this situation, the relation between the true redshift, $z_{\rm true}$, and wrong redshift, $z_{\rm misID}$, is
@@
If, as in the bottom panel, we select high-quality SPE galaxies (defined as having spe_z_proba $\geq\,0.99$ and a line width – expressed in terms of velocity dispersion – smaller than 680 km s-1 to exclude spurious broad features), we are left with three such features indicating line misidentifications. The one above the one-to-one diagonal corresponds to $(1+z_{\rm SPE})/(1+z_{\rm DESI})\simeq 1.76$: this results from cases in which SPE has interpreted a line as [O ii] when in fact it is $\mathrm{H}\,\alpha$. We can also notice two discrete features below the one-to-one relation, corresponding to $\mathrm{H}\,\alpha$ being misidentified as $\mathrm{Pa}\,\alpha$, $(1+z_{\rm SPE})/(1+z_{\rm DESI})\simeq 0.51$, or [S iii]9530, $(1+z_{\rm SPE})/(1+z_{\rm DESI})\simeq 0.68$. Finally, inspection of the $(1+z_{\rm SPE})/(1+z_{\rm DESI})$ histogram reveals the presence of a small peak at 1.34, corresponding to $\mathrm{H}\,\alpha$ being misidentified as $\mathrm{H}\,\beta$, although we surprisingly do not find the peak corresponding to [O iii].
As can be seen by comparing with Fig. 3, such misidentifications are largely suppressed by the $\mathrm{H}\,\alpha$ prior, and this aspect will be revisited in future versions of the pipeline. We remark that no obvious misidentification is observed in the target redshift range for cosmology, z=0.9z=0.9–1.8.
The second type of error (giving rise to what we call ‘noise interlopers’ in Risso et al., 2025) is caused by noise features or pipeline artefacts, such as unidentified residuals of the zeroth order signal, which are mistaken for real lines. In such cases, the measured redshift values are spread over a broad range. However, they are less likely to happen for redshift ranges where two real lines fall within the nominal NISP wavelength range, since this would require two spurious peaks separated by the same exact wavelength difference of two emission lines (this is the case in the range $1.4<z<1.8$ when considering [O iii]5007 and $\mathrm{H}\,\alpha$). Conversely, the $\mathrm{H}\,\alpha$ prior tends to push these spurious redshifts into the cosmology target range. As a result, the distribution of these incorrect redshifts is not uniform, as is shown in Fig. 4.
@@ 4.1 Broad categories
We consider two different cases. For the first, we selected high-quality spectra with three or more dithers on average (see Sect. 2.1). For the second case, we additionally required a class probability higher than 0.99. The results are not significantly better for the second case. This suggests that the spectral quality and the SPE classification probability criteria alone cannot yet efficiently select objects with a good classification.
The results are satisfactory for galaxies, with around 80% success. The results for quasars are poor, with about 80% of them being classified as galaxies. The current version of the pipeline uses a single template for quasars based on the average spectrum of SDSS quasars. It thus allows too little flexibility to fit the Euclid spectra reliably. In addition, this template has broad lines, and the observed type II AGNs tend to be fitted better by the galaxy templates, for example when the line-spread function is even slightly overestimated.
Only about 50% of the stars are correctly classified, while the remaining 50% leak into the galaxy class. Only 23% of these false galaxies are at $z<0.1$, and 21% of them are in the target redshift range for cosmology ($0.9<z_{\rm SPE}<1.8$). This emphasises the fact that additional non-SPE criteria will be necessary to minimise the contamination of the cosmological sample by stars and the high-quality imaging of the Euclid space telescope will undoubtedly be efficient in doing this. In contrast, about 75% of the objects classified as stars by SPE are DESI stars. The selection of star samples is thus very promising and the combination with other quality criteria from the Euclid photometric pipeline should provide high-quality samples.
Table 1: Confusion matrices of the OU-SPE classification.
@@
4.2 Galaxy sub-categories
In the spectro_zcatalog_spe_galaxy_candidates catalogue, the SPE PF provides a spe_subclass parameter, which is a short name for the emission line ratio templates used for the redshift calculation. The names refer either to templates based on morphological types (‘Irr’, ‘Scd’, ‘Sbc’) eventually coupled with additional properties derived from VVDS spectra such as ‘SB’ for starburst or ‘BX’ for excess in the blue. The most represented templates on the full sample or on the sample selected by $\texttt{spe\_z\_proba}\geq 0.99$ are either of starburst or ‘BX’ types. However, given the short wavelength range of NISP, the bias towards the solutions given by the $\mathrm{H}\,\alpha$ line and the fact that only fluxes greater than $2\times 10^{-16}\,\mathrm{erg\,s}^{-1}\,\mathrm{cm}^{-2}$ are detectable, this distribution is surely not representative of the reality of the observed population.
As for the continuum templates, the ones corresponding to recent single star formation episodes (5, 25, and 640 Myr) represent more than half of the sample. The fraction rises to 75% for the object with a correct redshift as compared to DESI. This confirms that the way the redshift is measured favours galaxies with high star formation rates.
@@ 5 Line-flux measurement products and performance
Once a redshift solution has been estimated for the galaxy model, the fluxes of all emission and absorption lines that are expected to be present in the spectrum are measured. It must be noted that for this Q1 release, even if an object is classified as a quasar or a star, the lines are identified at the redshift of the galaxy solution and therefore should not be trusted; this will be modified in future releases. Line fluxes are measured with two methods: ‘direct integration’ provides fluxes integrating the values of all the pixels of the emission line above the continuum; and ‘Gaussian fit’ derives them from the fit of the line using a single, double, or multiple Gaussian model, depending on the line considered. Given the low spectral resolution of NISP, the $\mathrm{H}\,\alpha$ and [N ii] doublet lines direct-integration fluxes are measured jointly over a unique wavelength interval to produce a single flux measurement, while the Gaussian fit provides fluxes separately for the $\mathrm{H}\,\alpha$ and [N ii] doublet. Examples of the fitting procedure results are displayed in Fig. 6 for a high-S/N emission line, as well as for one close to the flux detection limit.
Figure 6: Illustrative examples of the line flux measurement performed in SPE. Two spectra are shown from the EWS at $z\simeq 1.7$–1.8, with fluxes of $f(\mathrm{H}\,\alpha)\simeq 2\times 10^{-16}$ and $5.9\times 10^{-16}\,\mathrm{erg\,s}^{-1}\,\mathrm{cm}^{-2}$ (upper and lower panels, respectively). The left panels show the entire spectrum, highlighting the position of the main emission lines, and the right ones the results of the multiple Gaussian fit to the $\mathrm{H}\,\alpha$+[N $i$] complex showing in different colours the contribution of the $\mathrm{H}\,\alpha$ line (in blue), of the [N ii] doublet (in orange), and their combination (in red).
For each line and each method, SPE provides the flux, the S/N, the central wavelength, and the full width at half maximum. All measurements relative to emission lines are stored in the spectro_line_features_catalog_spe_line_features_cat table: for each object, this contains all the detected lines for each redshift solution. A given line can be identified either by its name (e.g. Halpha) or a numerical identifier, and the link between line names and identifiers is given in the spectro_model_catalog_spe_lines_catalog table. Therefore, to retrieve the flux of the $\mathrm{H}\,\alpha$ line for the main redshift solution of all objects, the selection should include spe_line_name==‘Halpha’ and spe_rank=0, together with any additional conditions.